A cash register contains only five dollar and ten dollar bills. It contains twice as many five’s as ten’s and the total amount of money in the cash register is 600 dollars. How many ten’s are in the cash register?
Answer by ikleyn(52786)

.

Group (mentally) all banknotes in sets, each containing 2 5-dollar bills and one 10-dollar bill.


Each group is worth  5+5+10 = 20 dollars.


Hence, there are 600/20 = 30 such groups.


Therefore, a cash register has 30 10-dollar bills and 2*30 = 60 5-dollar bills.   ANSWER

Solved (mentally).

-------------------

On coin problems,  see the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Solving coin problems mentally by grouping without using equations (*)
    - Non-typical coin problems
    - Santa Claus helps solving coin problem (*)
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

Specifically for grouping method, see the lesson marked (*) in the list.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.